### v1

@@ 2.1 Principle
2P-FOCUS corrects scattering in three steps: taking measurements, calculating the correction mask, and projecting the correction mask (Fig. 1a-c). First, 2P-FOCUS searches for open channels by projecting random intensity patterns in the Fourier domain while monitoring the fluorescence intensity excited under this frequency modulation (Fig. 1b, step 1). We refer to the ratio between the number of turned-on pixels and the total number of pixels as the sparsity of a random pattern. For example, a random pattern with 0.1 sparsity means that 10% of the pixels are turned on. The pixels on the DMD are binned into $n\times n$ pixel patches, termed “super-pixels”. Random patterns are generated using these super-pixels. Each pattern randomly selects multiple beams that may interfere at a focus after propagating through the unknown scattering media. If these beams interfere constructively, the center lobe of the interference pattern will excite fluorescence through two-photon absorption if the photon density in the center lobe is high enough. Otherwise, the beams do not generate a bright focus, resulting in low or zero fluorescence intensity. Beams that create a bright focus may also be multiply scattered, as long as their phase difference is zero. During our experiments, several thousand measurements are taken in this process. This process is highly nonlinear. The total fluorescence intensity is proportional to the fourth power of the number of interference beams (see Methods for the derivation).
The next step involves calculating the correction mask (Fig. 1b, step 2). The random masks that generate bright fluorescence are selected by thresholding the fluorescence intensity detected by the PMT. Since these beams can create bright center lobes through interference, they have little phase difference. Thus, summing these random patterns generates a new pattern. The multiple beams from this new pattern are coherent and in phase, producing a brighter center lobe. This new pattern serves as the correction mask, selecting beams corresponding to the open channels.
The final step is generating a binary correction mask for DMD projection by thresholding the intensity of the grayscale correction mask (Fig. 1b, step 3). To ensure a fair comparison, we maintain the same illumination power on the sample before and after correction, staying below the thermal damage threshold of biological tissues. For example, if a blank mask is used before correction, allowing no light blockage (Fig. 1a), we then increase the input power to the DMD when applying correction masks to compensate the power blocked by turned-off pixels (Fig. 1c). Power control can also been done by using different thresholds to binarize the grayscale correction mask. Different thresholds result in different illumination power on the sample given the same input power to the DMD, which will be discussed in detail in Section 2.3. As all beams selected by the grayscale correction mask are coherent and in phase, the subset of these beams after binarization remains coherent and in phase. Their interference produces a brighter center lobe compared to the scenario without scattering correction, not only effectively redistributing part of the illumination power from closed to open channels but also increasing the laser power in the center lobe due to coherent interference.
2.2 Experimental setup
The experimental setup is depicted in Fig. 1d and described in the Methods section. We utilized a femtosecond laser with a 1035 nm wavelength for two-photon excitation. Our setup included a near-infrared coated DMD with $1280\times 800$ pixels, of which only the central $800\times 800$ pixels were used, and a pattern rate of 12.5 kHz. Since the DMD is located on the conjugate plane of the back aperture, it induces spatial dispersion in the femtosecond pulses. To correct this spatial dispersion and compress the pulses, we implemented a reflective diffraction grating on the conjugate plane of the DMD [41, 42]. The parameters of the grating, the 4-f relay lenses (L3, L4), and the incident angles were carefully designed and aligned to fully compensate for the spatial dispersion. It’s important to note that the effective groove size of the DMD differs from the micromirror pitch size and was determined empirically. A two-axis scanning mirror system, operating at 1 kHz, is also positioned on the conjugate plane of the DMD. This system scans the focus across the field-of-view and currently represents the speed bottleneck of the system. The scanning mirrors are synchronized with the DMD, allowing the application of different correction patterns to specific sub-regions as necessary. This strategy significantly increases the system’s degrees of freedom, which is the product of the degrees of freedom of both the DMD and the scanning mirrors, thereby effectively controlling a 4D light field [43]. The modulated light field is relayed to the back aperture via 4-f systems and then focused by the objective lens onto the scattering sample. The sample is placed on a linear translation stage for capturing z-stack images. Two-photon excited fluorescence is collected by the objective lens and a collection lens, filtered through a color filter, and detected by a PMT.
2.3 Focusing through bone with 2P-FOCUS
Figure 2: Focusing through bone with 2P-FOCUS. (a) Representative random patterns with a sparsity of 0.4 and a super-pixel radius of 8 pixels. A total of 2,500 such random binary patterns are used for taking measurements of open channels. (b) Fluorescence intensity corresponding to the 2,500 patterns detected by the PMT is shown by the blue line. Random patterns contributing to the top 10% of fluorescence intensity are selected, as indicated by measurements above the red line. (c) The sum of the selected random binary patterns forms the grayscale correction mask. The grayscale intensity indicates the number of patterns contributing to each frequency patch. (d) A binary correction mask is generated by binarizing the grayscale correction mask. (e) Photo of the bone sample and the fluorescence sample. The bone sample is approximately 200-$\mu$m thick at the location where 2P-FOCUS is focused. The fluorescence sample is a microscope slide coated with a thin layer of fluorescent paint. (f) Zoomed-in image of the focus before correction, where no focus is identified. The peak intensity of the image is 10. (g) Zoomed-in image of the focus after correction, showing a clear focus with a peak value of 629. Images (f-g) were taken with a camera that replaced the PMT. (h) Comparison of the intensity profile along the x-axis before (blue line) and after correction (red line). (i) Fluorescence intensity before and after correction, as well as the intensity ratio, as functions of the laser power on the sample. (j) Fluorescence intensity before and after correction as a function of the sparsity of random patterns used for measurements. The radius of super-pixels is 8 pixels. (k) Fluorescence intensity before (blue) and after (red) correction as a function of the size of super-pixels. The sparsity of random patterns is 0.05. Fluorescence intensity before correction is measured with random patterns. (l) Fluorescence intensity before and after correction as a function of the number of measurements. The sparsity of random patterns is 0.01, and the radius of super-pixels is 12 pixels.
We first demonstrate 2P-FOCUS experimentally by focusing through a chicken bone onto a homogeneous fluorescence slide (Fig. 2). We adhered a piece of chicken bone without any exogenous fluorescent labels onto a microscope slide coated with a thin layer of red fluorescent paint (Fig. 2e). The surface of the bone is uneven and the part we focused through is about 200 $\mu$m. The bone is highly scattering and no focus is visible before correction (Fig. 2f) at an illumination power of 7 mW. The mean and standard deviation of intensity of Fig. 2f is $0.58\pm 1.01$. The maximum value of Fig. 2f comes from a hot pixel. In this experiment, We used random patterns with a sparsity of 0.4 and a super-pixel radius of 8 pixels. We will discuss about how sparsity and radius influence the result in the next paragraph. We increased the illumination power on the sample to 76 mW when we took measurements under random illumination, and decreased the illumination power back to 7 mW when we apply the correction mask. We projected 2500 random binary patterns (Fig. 2a) and recorded the corresponding fluorescence intensity (Fig. 2b). We then selected the random patterns corresponding to the top 10% brightest fluorescence (above the red line, Fig. 2b, 250 patterns) and summed them to form the grayscale correction mask (Fig. 2c). The grayscale value of the correction mask indicates the number of patterns that leave the super-pixel open. For instance, the peak value of 159 in Fig. 2c, located on the right bottom side in the Fourier domain, shows that most random patterns (159 out of 250) turned this pixel on to create a bright focus. We then adjusted the half-wave-plate to maximize the power input to the DMD, and binarized the grayscale correction mask (Fig. 2d) using a threshold that ensures the illumination power on the sample remains at 7 mW after modulation by the binary correction mask, which is the same as before correction. In this case, the correction mask only turns on pixles corresponding to ”completely opened” channels (see next paragraph for details). With this correction mask, we achieved a focus with a peak intensity of 629 (Fig. 2g). Figures 2f and 2g, captured by a camera replacing the PMT, display the focus’s image but were not used in the correction process. The bright focus after correction is clearly visible. The distorted shape of the fluorescence spot is due to the scattering of fluorescence as it propagates back through the bone. This occurs because our correction was aimed at maximizing total fluorescence intensity rather than preserving the shape. The intensity profile across the x-axis (Fig. 2h) also demonstrates the significant improvement in fluorescence intensity before and after correction.
We also explore how various factors influence the correction result, including the thresholds for binarization, the sparsity of random patterns, the radius of super-pixels, and the number of measurements. The binarization from Fig. 2c to Fig. 2d depends on the threshold applied to the grayscale correction mask, which controls the illumination power applied to the sample given the same input power to the DMD. The experiment depicted in Fig. 2a-h describes a scenario where the laser power on the sample is 7 mW, corresponding to the first data point in Fig. 2i. The fluorescence intensity value shown in Fig. 2i is measured by the PMT, so it may differ from the values in Figs. f-h. As we increase the illumination power (to 15 mW, 34 mW, 52 mW, and 76 mW), we observe that the fluorescence intensity before correction increases quadratically with the illumination power (see Extended Data Fig. S1a), consistent with the principles of two-photon excitation. Interestingly, the fluorescence intensity after correction, achieved by applying different thresholds to Fig. 2c (see Extended Data Fig. S1b-c), also increases, but does not follow the quadratic rule. This deviation is attributed to the effect of open channel coupling. The brightest pixels in Fig. 2c represent ”completely opened” channels, while other bright pixels indicate ”partially opened” channels. Activating the DMD pixels corresponding to the partially opened channels increases the illumination power on the scattering sample. However, only a portion of this power is delivered to the fluorescence slide and constructively interferes with other beams. The most significant improvement in fluorescence intensity is achieved when the laser power on the sample is low. In other words, the strategy to achieve a large improvement after correction is making the correction mask turn on only the pixels corresponding to ”completely open” channels, while ensuring that the power after modulation is still high enough to achieve a sufficient signal-to-noise ratio. Therefore, in Fig. 2a-h, we chose to maximize the laser power input to the DMD and set a high threshold for binarization in order to couple light into completely open channels and achieve the largest improvement. Fig. 2i clearly demonstrates the presence of open channels in the scattering bone.
The second factor we explore is the sparsity of random patterns, which indicates the percentage of pixels that are turned on out of the total number. For random patterns with a super-pixel radius of 8 pixels, there are 2,500 super pixels in each mask, and a total of 2,500 masks are used, forming a $2500\times 2500$ matrix $P$. For sparsity ranging from 0.05 to 0.9 (see an example in the Extended Data Fig. S2a), these matrices are all orthogonal, as $P^{T}P=I$. The sparsity of random patterns is also related to illumination power. With the same input laser power to the DMD, less sparse patterns block less input power and emit more power to the sample. This increases the signal-to-noise ratio of the raw measurements, consequently improving the accuracy of the correction masks. In the case shown in Fig. 2j, we first measure the fluorescence intensity using random patterns at 7 mW, as in the pre-correction case, and then take raw measurements under different sparsity (0.05, 0.1, 0.2, 0.3, and 0.4), while keeping the input power to the DMD constant. These measurements generate five different correction masks, each with an output power of 7 mW (see Extended Data Fig. S2b). We also took raw measurements under sparsity ranging from 0.2 to 0.5 through another region of the bone and found that the fluorescence signal decreases at the sparsity of 0.5 (see Extended Data Fig. S2c-d). The correction mask derived from the random patterns with 0.4 sparsity shows the most significant improvement in fluorescence intensity in both cases, as depicted in Fig. 2j and Fig.S2c. The random patterns with a sparsity of 0.4, along with the corresponding measurements and the correction mask, are shown in Fig. 2a-d. In practice, the optimal sparsity of random patterns depends on the scattering properties of the sample, and the brightness of the fluorescence signal, and the maximum power of the laser.
We also explored how the radius of super-pixels influences the scattering correction. We used the center square region of the DMD, consisting of $800\times 800=640,000$ pixels. Well-sampling all these pixels requires 640k random patterns, which is too large for the on-chip RAM of the DMD to handle and also requires more computing and experimental time. Therefore, we binned these pixels into super-pixels to accelerate the process. For super-pixels with a radius of $w$, the number of random patterns to well-sample the frequency domain is $800^{2}/(2w+1)^{2}$. We compared the results generated by patterns with $w$ = 8, 12, and 21 (see Fig. 2k). The fluorescence intensity before correction was measured under random pattern modulation with a sparsity of 0.01. The illumination power on the sample was approximately 12 mW, similar before and after correction for all three pattern types. The results indicate that patterns with smaller super-pixels correct scattering more effectively because they identify open channels with greater precision.
We also explored how the number of measurements influences scattering correction. For random patterns with a super-pixel radius of 12 pixels, the minimum number of measurements required for well-sampling is 1024. We generated correction masks with various numbers of raw measurements, ranging from 200 to 2000, and compared the results of the scattering correction. As shown in Fig. 2i, the results improve as the number of measurements increases. However, the improvement becomes trivial with 1024 or more measurements. Considering the computing and experimental time, the results suggest that we should take the minimum number of measurements that adequately sample the Fourier domain, without oversampling.
2.4 Imaging deep in the mouse brain with 2P-FOCUS
Figure 3: Imaging fluorescence-labeled neurons deep in the mouse brain using 2P-FOCUS. (a) The volumetric view of parvalbumin (PV) interneurons expressing cell-fill tdTomato imaged by 2P-FOCUS in vitro. The image volume is $230\times 230\times 450\mu m^{3}$. (b) The top plane of the image stack shows the dendrites and axons of PV interneurons near the surface of the mouse brain, with no correction applied. (c) The bottom plane of the image stack, 450 $\mu$m deep from (b), shows some cell bodies. Sub-region correction is applied here, and the illumination power is increased compared to (b). The background signal originates from the dendrites and axons of PV interneurons, as cell-fill expression, rather than sparse expression, is used in the experiment. (d) Schematic diagram illustrating global correction and the correction mask used for (e-f). The same mask is used for all scanning locations. The correction mask is generated by optimizing the fluorescence intensity of the reference neuron, indicated by the yellow arrow in (e). (e-f) Images of the same region at 250 $\mu$m depth (e) before and (f) after global correction. Fluorescence intensity is improved by about 3-fold after correction. (g) Schematic diagram for sub-region correction and the sub-regions in (h-i). (h-i) Images of the same region at 400 $\mu$m depth (h) before and (i) after sub-region correction. Fluorescence intensity is improved by about 9.3-fold after correction. (j) The correction masks for four sub-regions used to capture image (i). (k) Comparison of the intensity profile of representative neurons at 250 $\mu$m depth before and after correction (indicated by the blue arrow in (e) and the red arrow in (f)). (l) Comparison of the intensity profile of representative neurons at 400 $\mu$m depth before and after correction (indicated by the blue arrow in (h) and the red arrow in (i)).
2.4.1 Imaging neurons deep in the brain with global and sub-region scattering correction.
We next demonstrate the imaging capabilities of 2P-FOCUS by imaging fluorescence-labeled neurons in the mouse brain. 2P-FOCUS can actively correct scattering during the imaging process with point scanning; that is, it scans the focus across the field-of-view pixel by pixel while applying corresponding correction masks. In our first imaging experiment, we imaged a $230\times 230\times 450\mu m^{3}$ volume in the primary visual cortex of a transgenic mouse brain with cell-fill expression of tdTomato in all parvalbumin (PV) interneurons (see Fig. 3a). The mouse underwent transcardial perfusion with 4% PFA and was then preserved in 4% PFA until imaging. During the imaging session, the whole brain was maintained in phosphate-buffered saline and placed beneath the objective lens. The brain was imaged from the top downward, starting from a plane near the surface. Without any correction, the top plane shows the dendrites and axons of PV interneurons (see Fig. 3b). When imaging the top 250 $\mu$m, we did not apply any correction but gradually increased the illumination power on the sample from 2 mW to 14 mW. The first correction was applied when imaging the plane at a 250 $\mu$m depth (see Fig. 3d-f). Before correction, the image contrast is low, but we can still identify some bright neurons (Fig. 3e). We selected one of the bright neurons near the center of the field-of-view, indicated by the yellow arrow in Fig. 3e, as the reference object. It is important to note that the reference neuron is not a guide star, as it does not need to be a point source; any fluorescent object can be used as a reference. We applied an offset voltage to the scanning mirrors to focus on the reference neuron, and then performed a scattering correction to maximize the fluorescence intensity at this neuron. We used 2,500 random masks with a super-pixel radius of 8 pixels and a sparsity of 0.4 for measurements, and generated a correction mask as shown in the bottom of Fig. 3d. We applied this correction mask to all scanning positions, referring to as a global correction, and acquired the image of the same region after correction (Fig. 3f). The maximum fluorescence intensity is improved by about 3 folds (from 27 to 80). The image contrast is also improved after correction. For example, the three neurons pointed by the arrow (Fig. 3e-f) were not distinguishable before correction but became distinguishable after correction. The intensity profile of the three neurons are plotted in Fig. 3k.
As the imaging depth increases, the global correction mask no longer provides high intensity and high contrast images when the imaging region is larger than the range of the memory effect. We performed corrections at 290 $\mu$m and 380 $\mu$m depths using sub-region correction (see Fig. 3g-j). Sub-region correction involves selecting multiple reference points located in different areas to generate a correction mask for each sub-region, rather than using a single mask for the entire field-of-view. The scanning mirrors are synchronized with the DMD, allowing each sub-region correction mask to be applied when scanning the corresponding regions. Figs. 3h-i show the image at 400 $\mu$m depth before and after sub-region correction. before and after sub-region correction. Four correction masks are generated using reference fluorescent objects located at 380 $\mu$m depth (see Fig. 3g, 3j, and Supplementary Figure X). We initially split the entire field-of-view into 3x3 sub-regions; however, only four out of the nine sub-regions contained bright objects that could serve as references. Therefore, we generated four correction masks, each corresponding to one of these references. We then merged the nearest sub-regions without references into those with references to form four new sub-regions (see Fig. 3g, bottom). We applied these correction masks to the plane at 400 $\mu$m depth, which differs from the correction plane but is within the range of memory effect, and successfully improved both the fluorescence intensity and image contrast (see Fig. 3i) compared to the pre-correction image (Fig. 3h).The intensity profile of a representative neuron, shown in Fig. 3l, indicates that the fluorescence intensity improved by approximately 9.3 times. We noticed that the improvement is less than that achieved when focusing through bone. This is because the fluorescence paint used in the focusing experiment is much brighter than the fluorescence-labeled neurons in the mouse brain. With brighter fluorophores, the correction mask can selectively turn on only the pixels corresponding to completely opened channels and still achieve a sufficient signal-to-noise ratio. In contrast, with less bright fluorophores, the correction mask has to turn on more pixels, including those corresponding to ”partially opened” channels. Therefore, the intensity improvement in brain imaging is less than that in focusing through bone, which is consistent with our result in Fig. 2i. Sub-region correction enables us to image a $230\times 230\mu m$ region deep within the scattering tissue. Therefore, for highly scattering media and large fields-of-view, 2P-FOCUS with sub-region correction can correct scattering outside of an isoplanatic patch, achieving better results than global correction.
2.4.2 Imaging cerebral blood vessels deep in the mouse brain by precise open-channel coupling.
To further improve the correction and demonstrate the broad applicability of 2P-FOCUS, we imaged another mouse brain in which cerebral blood vessels were labeled through intravascular FITC-dextran injection, using patterns with a super-pixel radius of 4 pixels. Similar to the setup in Fig. 3, the mouse brain was immersed in phosphate-buffered saline and imaged from the top downward. The top plane that we imaged was not on the brain’s surface but below it, as indicated by the visible cross-sections of cerebral blood vessels. We imaged the top 270 $\mu$m volume without correction, meaning all pixels of the DMD were turned on. The maximum intensity projection of the top 100 $\mu$m-thick volume is shown in Fig. 4b, where the cerebral blood vessels are clearly visible. At the bottom of the volume (410 $\mu$m - 510 $\mu$m depth), blood vessels remain visible after scattering correction (see Fig. 4c). Since blood vessels span three-dimensionally, we chose to show the maximum intensity projection rather than a single plane. As the imaging depth increased, we gradually raised the laser power on the sample, from 3 mW (for $Z=0-160\mu m$) to 19 mW (for $Z=250-510\mu m$). We first performed scattering correction at a depth of 270 $\mu$m in three sub-regions, using 2,500 random patterns with a super-pixel radius of 8 pixels in each sub-region. At a depth of 300 $\mu$m, we conducted another sub-region correction with 4 sub-regions, using 8,100 random patterns with a super-pixel radius of 4 pixels in each sub-region (Extended Data Fig. S4a-c). The same 4 correction masks (see Fig. 4h-i) were used when imaging at a depth of 340 $\mu$m (Fig. 4e), revealing fine structures of cerebral capillaries in high contrast (refer to the zoomed-in view in Fig. 4g). Compared to the uncorrected image with the same illumination power (Figs. 4d, 4f), where all pixels on the DMD were turned on, these capillaries were barely visible. We compared the fluorescence intensity of a cross-section of these capillaries in the images before and after correction (Fig. 4j), indicating a 30.6-fold improvement in peak intensity after correction. We also performed another correction at 390 $\mu$m depth with six sub-regions as the imaging depth increases (see Extended Data Fig. S4d-f). Compared to the results in Fig. 3, the outcomes in Fig. 4 show better scattering correction, attributable to the use of patterns with smaller super-pixel radii that generate more precise correction masks. Additionally, the fluorescence labeling of blood vessels is brighter than that of the neurons, despite 1035 nm not being the optimal two-photon excitation wavelength for FITC-dextran. Therefore, the signal-to-noise ratio of measurements under random pattern modulation is higher, resulting in a more accurate correction mask for more precise open-channel coupling. This result demonstrates successful coupling of light into open channels, significantly enhancing laser light propagation through the scattering brain tissue.
Figure 4: Imaging blood vessels with intravascular fluorophore injection deep in the mouse brain using 2P-FOCUS. (a) Volumetric view of cerebral blood vessels with intravascular FITC-dextran injection, imaged by 2P-FOCUS in vitro. The image volume is $230\times 230\times 510\mu m^{3}$. (b) Maximum intensity projection (MIP) of the top 100 $\mu$m-thick volume along the z axis. (c) MIP of the bottom 100 $\mu$m-thick volume along the z axis. (d) Two-photon image of blood vessels at 340 $\mu$m depth without correction. All pixels of the DMD are turned on. The illumination power on the sample is 19 mW, with a peak fluorescence intensity of 27. (e) Image of the same region as in (d) after sub-region correction. The illumination power on the sample is the same as in (d) at 19 mW, but the peak fluorescence intensity is improved by 30.6-fold, reaching 826. (f-g) Zoomed-in view of the regions in the dashed box in (d-e), showing that small capillaries become visible after scattering correction. (h) The location of the 4 sub-regions on the image plane. (i) The correction masks used in the experiment when acquiring image (e). The radius of super-pixels is 4 pixels and the sparsity is 0.4. (j) Comparison of the fluorescence intensity profile before and after correction along the dashed line in (f-g). The intensity profile of the case before correction is magnified by 10 times for display purposes.
3 Discussion
We have demonstrated that 2P-FOCUS is a novel method for correcting scattering in two-photon microscopy. 2P-FOCUS identifies and couples femtosecond laser light into open channels within scattering tissue by modulating the intensity of light in the Fourier domain, that is, by controlling multiple beam interferences. We demonstrated 2P-FOCUS by focusing through chicken bone, where the focus was initially invisible became distinct after correction. We quantitatively evaluated how various factors affect the correction result, including the number of open channels, the sparsity of random patterns, the super-pixel radius of the patterns, and the number of measurements. We also demonstrated the imaging capability of 2P-FOCUS by capturing image stacks of fluorescence-labeled neurons and cerebral blood vessels in a mouse brain in vitro. The intensity and signal-to-noise ratio of these images were significantly improved compared to images without correction (where all pixels on the DMD were turned on), enabling access to regions over 500 $\mu$m deep in the mouse brain. With sub-region correction, we could correct scattering outside the range of the memory effect, enabling imaging across a $230\times 230\mu m^{2}$ field-of-view deep in the mouse brain. Therefore, 2P-FOCUS can effectively correct scattering and enhance fluorescence signals in a large volume.
2P-FOCUS could be improved in two key aspects. First, to extend the imaging depth further, 2P-FOCUS needs to acquire high signal-to-noise ratio measurements even deep inside the sample. This proves challenging for the current version of 2P-FOCUS, as measurements are taken under completely random modulation in the Fourier domain. This limitation could potentially be overcome by incorporating the correction mask from the plane above to generate the correction mask for the current plane. Second, the current process of generating a correction mask in 2P-FOCUS is not fast enough to correct millisecond-scale dynamic scattering due to blood flow [44]. For instance, collecting 2,500 measurements at a 1 kHz sampling rate takes 2.5 seconds. The computational processes of summation and binarization take about 0.5 seconds, and projecting the correction mask onto the DMD takes 1 millisecond. Consequently, the total time to generate one correction mask is approximately 3 seconds. The speed of this process could be further enhanced with a higher bandwidth data acquisition card, faster scanning mirrors, a higher repetition rate laser, and software designed for auto-correction with closed-loop control.
In summary, 2P-FOCUS is a novel two-photon microscopy system with enhanced scattering control. It can actively boost fluorescence intensity by a few tens of times at the same illumination power on the sample. This capability makes it a powerful tool for deep tissue imaging, facilitating studies in neuroscience, immunology, and cancer.
4 Methods
4.1 2P-FOCUS theory
We describe the process of 2P-FOCUS, which is based on multiple beam interference through scattering media. In the first step, the DMD modulates a collimated laser beam using binary random patterns on a plane relayed to the objective lens’s back aperture. If N super-pixels with a radius of $W$ are turned on, the electric field after intensity modulation is
@@ 4 Methods
where $rect$ is the rectangular function describing a turned-on super-pixel. The next step is to calculate the electric field $E_{0}$ on the surface of scattering medium (assuming the uneven of surface is negligible compared to the working distance of objective lens), right before it propagates into the scattering medium. If the thickness of the scattering medium is $\Delta z_{0}$, $E_{0}$ is calculated by
@@
That is, to calculate the incident field on the surface of the sample, we first calculate the electric field at the focus and then back propagate it from the focus to the surface over a distance of $\Delta z_{0}$. Combining Eq. 2 and Eq. 3, the electric field on the surface of the scattering medium can be simplified to
$E_{0}(x,y)=\mathcal{F}^{-1}{exp[-j2\pi\Delta z_{0}\sqrt{1/\lambda^{2}-(\xi^{2}+\eta^{2})}]\cdot E_{A}(\xi,\eta)}.$ (4)
@@
where $H$ is the forward Fresnel propagation over distance $\Delta z_{0}$. The incident electric field $E_{0}$ propagates in the scattering medium over distance $\Delta z_{0}$, reaching the reference fluorescence object. The excitation field on the fluorescence object is
@@
where $T_{exc}$ is the transmission matrix of the scattering medium. Notice that 2P-FOCUS does not calculate the transmission matrix. The intensity of the excitation field $E_{exc}$ is
@@
Since fluorescence intensity is quadratically proportional to the intensity of excitation light. Therefore, fluorescence intensity excited by two-photon process is
@@
The excited fluorescence propagates backward to exit the scattering medium. Because the wavelength of te fluorescence is different from that of the excitation light, and fluorescence is incoherent (or partial coherent) light while excitation light is coherent light, we use an intensity transmission matrix $T_{f}$ for fluorescence. This matrix is different from the transmission matrix for excitation light, $T_{exc}$. The signal detected by the PMT is the total fluorescence intensity exiting the scattering medium:

### v2

@@ 2.1 Principle
Figure 1: Principle and schematic diagram of 2P-FOCUS. (A) Without correction, all pixels of the digital micromirror device (DMD) are turned on, the incident light is scattered and cannot form a tight focus. (B) The process of generating a correction mask involves three steps. (C) With correction, the binary correction mask is projected on the DMD, allowing only the beams that interfere constructively pass through. The illumination power on the sample is maintained the same before and after correction by increasing the input power to the DMD to compensate for the power loss due to turning off some pixels. The correction results in a brighter focus compared to the case before correction. (D) Optical schematic diagram of 2P-FOCUS. Details are in Appendix A. (E) The lateral point-spread-function (PSF) before (left) and after (right) dispersion compensation, measured by a camera in reflection mode. (F) The lateral (left) and axial (right) PSF without scattering media, measured by the PMT in (D). (G) Intensity profile of the PSFs in (F). The lateral resolution is 1.7 $\mu$m and the axial resolution is 6.6 $\mu$m without scattering media.
2P-FOCUS corrects scattering in three steps: taking measurements, calculating the correction mask, and projecting the correction mask (Fig. 1A-C). First, 2P-FOCUS projects random intensity patterns in the Fourier domain while monitoring the fluorescence intensity excited under this modulation (Fig. 1B, step 1). In the following experiments, random patterns are generated from a uniform discrete distribution on [0, 1] using MATLAB, with the random seed shuffled each time a new set of random patterns is created. Different sets of random patterns, consisting of hundreds to thousands of patterns, are generated for varying sparsity levels and super-pixel sizes. Random intensity patterns has been used as an orthogonal basis in computational optics [41]. Random patterns are orthogonal to each other because each of them consists of a large number of segments (hundreds to thousands), thus satisfying the mathematical criteria of an orthogonal matrix, as $P^{T}P=I$, where $P$ is the reshaped and normalized random pattern matrix and $I$ is the identity matrix. In other words, the dot product of any two distinct random patterns is approximately zero. Each segment on the DMD consists of $n\times n$ pixels, termed “super-pixels”. Random patterns are generated using these super-pixels. If $N$ super-pixels with a radius of $W$ are turned on, the electric field after intensity modulation is
@@
where $rect$ is the rectangular function describing a turned-on super-pixel. The next step is to calculate the electric field $E_{0}$ on the surface of scattering medium, assuming the unevenness of the surface within the beam size is negligible compared to the working distance of the objective lens. If the thickness of the scattering medium is $\Delta z_{0}$, $E_{0}$ is calculated by
@@
That is, the incident field on the surface of the sample is calculated by first computing the electric field at the focus, and then back-propagating it from the focus to the surface over a distance of $\Delta z_{0}$. By combining Eq. 2 and Eq. 3, the electric field on the surface of the scattering medium can be simplified to
$E_{0}(x,y)=\mathcal{F}^{-1}{\exp[-j2\pi\Delta z_{0}\sqrt{1/\lambda^{2}-(\xi^{2}+\eta^{2})}]\cdot E_{A}(\xi,\eta)}.$ (4)
@@
where $H$ is the forward Fresnel propagation over distance $\Delta z_{0}$. The incident electric field $E_{0}$ propagates through the scattering medium over a distance $\Delta z_{0}$, eventually reaching the reference fluorescence object. The excitation field on the fluorescence object is
@@
where $T_{exc}$ is the transmission matrix of the scattering medium at the excitation wavelength. The intensity of the excitation field $E_{exc}$ is
@@
Since each pattern randomly selects multiple beams, they may or may not constructively interfere at a focus after propagating through the unknown scattering media. If these beams interfere constructively, the center lobe of the interference pattern will excite fluorescence through two-photon absorption if the photon density in the center lobe is high enough. Otherwise, the beams do not generate a bright focus, resulting in low or zero fluorescence intensity. Beams that create a bright focus may also be multiply scattered, as long as their phase difference is zero. In two-photon microscopy, fluorescence intensity is quadratically proportional to the intensity of excitation light in two-photon excitation, represented as
@@
The excited fluorescence propagates backward to exit the scattering medium. Because the wavelength of the fluorescence differs from that of the excitation light, and because fluorescence is incoherent (or partial coherent) light while the excitation light is coherent, we use an intensity transmission matrix $T_{f}$ for fluorescence. This matrix differs from the transmission matrix $T_{exc}$ used for excitation light. The signal detected by the PMT represents the total fluorescence intensity exiting the scattering medium, and is expressed as: